OpenStudy (anonymous):

A sum of $10 000 is placed in a bank account and earns 12% interest per annum, compunded annually. How much money is in the account at the end of 6 years, just after the final interest has been paid?

OpenStudy (apoorvk):

10,000 is the initial money, and then next year I have 112% of it, right?

OpenStudy (apoorvk):

And every year, I have 112% of the previous year's amount, right?

OpenStudy (anonymous):

This is what I did. Let A_n be the amount at the end of nth year. \[A_1 = 10 000\left(1+\frac{12}{100}\right) ^{1} = 10 000(1.12)\] \[A_2 = 10 000(1.12)+10 000)(1.12) => 10 000(1.12)^2+ 10 000(1.12) \] . . \[A_6 = 10 000(1.12)^6+10 000(1.12)^5+....+10 000(1.12)\] \[= 10 000(1.12)+.....+10 000(1.12)^5+10 000(1.12)^6 \] then its a G.P \[\frac{10000(1.12)(1.12^{6}1)}{1.12-1} \] but apparently the answer is wrong.. i havent done these in a long time im not sure..

OpenStudy (apoorvk):

So, at the end of 'n', years, I have my original amount made 112% 'n' times, over and over again. So, I have at the end of 'n' years, amount 'A', which is \(A = 10000(1.12)^n\) Oh no, you just need the nth term not the sum!

OpenStudy (apoorvk):

The bank, unfortunately, till now has no such investment-programs, where we receive the sum of our account balances at the end of each year.

OpenStudy (anonymous):

it says at the end of 6 years; then the whole thing right? lol

OpenStudy (anonymous):

nevermind i figured it out

OpenStudy (apoorvk):

good then..